Solve the equation in exact differentials: (3x2+2y)dx+(2x-3)dy=0


1
Expert's answer
2021-11-29T15:51:52-0500

"(3x^2+2y)_y=(2x-3)_x=2"


"F=\int (3x^2+2y)dx=x^3+2xy+g(y)"


"F_y=2x+g'(y)=2x-3"

"g'(y)=-3"

"g(y)=-3y+c"


"F=x^3+2xy-3y+c"


"x^3+2xy-3y+c=0"